Fermi National Laboratory


Accelerator Update for September 14, 2005 - September 16, 2005

Wednesday September 14
The day shift began with Operations monitoring store 4386 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Operations began preparing the Recycler to stash antiprotons at 8:53 AM.

Operations reported at 9:18 AM that the Recycler stash was on hold while experts investigated a problem with a correction element.

The Recycler stash ended and Pbar resumed stacking at 10:01 AM.

At 11:02 AM, the H- Source tripped off due to its generator brushes failing. Experts accessed the Preaccelerator to change the brushes.

The Preaccelerator worked ended at 11:31 AM and Operations reestablished beam to all areas.

Operations terminated store 4386 at 8 PM.

Operations put the Tevatron (TeV) into shot setup at 8:13 PM.

Operations established combination store 4387 at 11:19 PM with an initial luminosity of 67.99E30. Pbar and MCenter resumed normal operations.

Thursday September 15
The midnight shift began with Operations monitoring store 4387 and a small stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began stashing antiprotons at 3:36 AM and finished at 4:23 AM. Pbar and MCenter resumed normal operations.

The Recycler began stashing at 9:01 AM and finished at 9:53 AM. Pbar resumed stacking.

MCenter resumed taking beam at 10:46 AM.

The Recycler began stashing at 2:18 PM.

At 2:42 PM, Operations reported that Pbar lost 12 mA of their stack just before the last transfer of antiprotons to Recycler. Investigations showed that a dipole shunt (A:BS410) dropped to zero amps. A controls expert replaced a 054 card.

At 3:35 PM, Operations reported that the stash had ended, but that Pbar had resumed stacking.

MCenter resumed taking beam at 3:58 PM.

Pbar resumed stacking a 4:01 PM.

Main Injector (MI) experts began investigating beam problems at 5 PM.

Friday September 16
The midnight shift began with Operations monitoring store 4387 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4387 at 12:07 AM.

Operations put the TeV into shot setup at 12:29 AM.

Operations established combination store 4388 at 3:19 AM with an initial luminosity of 108.3E30. Pbar and MCenter resumed normal operations.

The Recycler began stashing at 6:14 AM and ended at 7:20 AM. Pbar and MCenter resumed normal operations.

At 7:17 AM, Operations reported the H- Source high voltage was sparking. Operators investigated and replaced a regulator module.

The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/16/2005   email Fermilab